OpenStudy (anonymous):

solve by completing the square. x squared +14x=15. show your work :)

OpenStudy (lgbasallote):

x^2 + 14x = 15 x^2 + 14x + (14/2)^2 = 15 + (14/2)^2 x^2 + 14x + 7^2 = 15 + 7^2 x^2 + 14x + 49 = 15+49 (x+7)^2 = 64 x+7 = sqrt (64) x + 7 = +/- 8 x = -7 +/- 8 x = 1 x = -15